A mixture of noble gases consisting of 5.50g He and 15.0g Ne was placed in a cylinder at 1atm.What is the partial pressure of Helium in the mixture?


1
Expert's answer
2021-03-29T06:09:56-0400

Molar Mass of He = 4.003

Molar Mass of Ne = 20.1797

= 5.5/4.003 = 1.365 moles of He

= 15/20.1797 = 0.743 moles of Ne

Total moles = 1.365 + 0.743 = 2.108

P He = 1.365/2.108 × 1

= 0.648 atm